A new start-up company promises to pay an investor each quarter for the next two years. The company will pay $20,675.00 per quarter for the first four quarters, and then $25,775.00 per quarter for the following four quarters. If the investor wants a 11.68% APR return with quarterly compounding, what is the value of the investment opportunity today?
Answer #1
1) The quarterly interest rate = 11.68%/4 = 2.92%
2) PV of the payment for the first 4 quarters = 20675*(1.0292^4-1)/(0.0292*1.0292^4) = $     76,998.26
3) Discounted value of the second 4 quarterly payments at EOY1, that is at the end of the 4th quarter = 25775*(1.0292^4-1)/(0.0292*1.0292^4) = $      95,991.78
PV of the second stream of quarterly payments = 95991.78/1.0292^4 = $     85,552.94
Value of the investment opportunity $ 1,62,551.20